OpenStudy (anonymous):

In this circuit, which resistor will draw the least power? will draw inside!! A. R_1 B. R_2 C. R_3 D. the power will be the same for each resistor

OpenStudy (anonymous):



OpenStudy (michele_laino):

here the current is: I=12/(2.2+1.1+2.7)=...amperes

OpenStudy (anonymous):

2!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the power dissipated by R_1 is: P_1= R_1 * I^2= 2.2* 4=...

OpenStudy (anonymous):

8.8!

OpenStudy (michele_laino):

ok! the power dissipated by R_2, is: P_2= R_2 I^2 = 1.1* 4=...watts

OpenStudy (anonymous):

4.4!

OpenStudy (michele_laino):

ok! finally the power dissipated by R_3, is: P_3 = R_e*I^2 = 2.7*4=...watts

OpenStudy (michele_laino):

oops.. P_3= R_3* I^2

OpenStudy (anonymous):

10.8!

OpenStudy (michele_laino):

ok! reassuming, we have these dissipated powers: P_1=8.8 watts, P_2=4.4 watts P_3=10.8 watts which is the littlest value among them?

OpenStudy (anonymous):

P2! so is choice B our solution?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay!! thanks!:)

OpenStudy (michele_laino):

:)